# Integrated soliton microcombs beyond the turnkey limit

### v1

@@ Results
Experimentally, we operate in the extended detuning regime with AR and NR heater powers of 127 mW and 1 mW, respectively. Tuning the AR yields an approximately linear decrease in the repetition rate of 63.9-63.9 kHz/mW and a concomitant increase in the pump frequency of 6.1 MHz/mW (Fig. 4a). Near this operating point, tuning only the NR produces a tenfold smaller change in repetition rate (6-6 kHz/mW) but a much larger pump-frequency shift of 26.2 MHz/mW (Fig. 4b). Based on these response coefficients, we implement orthogonal control by tuning the AR and NR in opposite directions with an appropriate $\sim 5{:}1$ ratio of heater-power changes (Fig. 4c). This strategy yields an effective repetition-rate tuning coefficient of 47.3-47.3 kHz/mW (normalized to $|\Delta P_{\mathrm{AR}}|+|\Delta P_{\mathrm{NR}}|$), while suppressing the pump-frequency tuning to only 0.35 MHz/mW. Although the orthogonality can be further refined to reduce residual cross-talk, the demonstrated control is already sufficient to enable potential full stabilization of the soliton microcomb 38, 39.

### v2

@@ Results
Experimentally, we operate in the extended detuning regime with AR and NR heater powers of 127 mW and 1 mW, respectively. Tuning the AR yields an approximately linear decrease in the repetition rate of 63.9-63.9 kHz/mW and a concomitant increase in the pump frequency of 60.8 MHz/mW (Fig. 4a). Near this operating point, tuning only the NR produces a tenfold smaller change in repetition rate (6-6 kHz/mW) but a much larger pump-frequency shift of 261.7 MHz/mW (Fig. 4b). Based on these response coefficients, we implement orthogonal control by tuning the AR and NR in opposite directions with an appropriate $\sim 5{:}1$ ratio of heater-power changes (Fig. 4c). This strategy yields an effective repetition-rate tuning coefficient of 47.3-47.3 kHz/mW (normalized to $|\Delta P_{\mathrm{AR}}|+|\Delta P_{\mathrm{NR}}|$), while suppressing the pump-frequency tuning to only 3.5 MHz/mW. Although the orthogonality can be further refined to reduce residual cross-talk, the demonstrated control is already sufficient to enable potential full stabilization of the soliton microcomb 38, 39.